OpenStudy (anonymous):

how do you find the sum of this: \[\sum_{n=0}^{15}2\left(\begin{matrix}4 \\ 3\end{matrix}\right)^{n}\]

OpenStudy (kinggeorge):

Find \[\binom{4}{3}\]In this case, 4. Then use the expansion for a finite geometric series to find what the sum of \(4^n\) from 0 to 15 is. Finally multiply by 2.

OpenStudy (anonymous):

can you show me the steps?

OpenStudy (kinggeorge):

\[\binom{4}{3}=4\] The sum of a finite geometric series of \(x^n\) from \(n=0\) to \(n=k\) is \[x^{k+1}-1 \over x-1 \]Thus, your sum should be \[2 \cdot \left({4^{16}-1 \over 4-1} \right)=2\cdot\left({4^{16}-1 \over 3}\right)\]You could calculate the actual number this gave you if you wanted to.

OpenStudy (anonymous):

so my answer woul be 28263311530?

OpenStudy (kinggeorge):

I'm getting 2863311530. I think you put an extra 2 in there.